if five pencils and three erasers cost 65 cents, and three pencils and two erasers cost 40 cents, what is the cost of one pencil and one eraser
Answer by lwsshak3(11628)

if five pencils and three erasers cost 65 cents, and three pencils and two erasers cost 40 cents, what is the cost of one pencil and one eraser.
**
let x=cost of one pencil
let y=cost of one eraser
..
5x+3y=65
3x+2y=40
..
10x+6y=130
9x+6y=120
subtract
x=10
..
2y=40-3x
2y=40-30=10
y=5
..
cost of one pencil=10 cents
cost of one eraser=5 cents